OpenStudy (darkigloo):

How to separate the variables and find the integral of dy/dx=y^2/x^3

OpenStudy (darkigloo):

This is what I have so far: \[x^3dy=y^2dx\] \[\int\limits(dy/y^2)=\int\limits(dx/x^3)\]

OpenStudy (anonymous):

\[\int\limits y ^{-2}dy=\int\limits x ^{-3}dx+c\] solve

OpenStudy (darkigloo):

how did you get that? please show step by step

OpenStudy (anonymous):

\[\frac{ 1 }{ x ^{m} }=1*x ^{-m}\] it is the next step to what you have written.

OpenStudy (darkigloo):

oh! ok. so if i do the integral i'll get \[\frac{ 1 }{ y }=-\frac{ 1 }{ 2} x ^{-2 } +C\] is that correct?

OpenStudy (darkigloo):

sorry, \[-\frac{ 1 }{ y }\]

OpenStudy (anonymous):

correct

OpenStudy (darkigloo):

Thank you!